OpenStudy (anonymous):

Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

point slope form

OpenStudy (anonymous):

y-y0 = m(x-x0)

OpenStudy (anonymous):

that doesnt help me

OpenStudy (anonymous):

that's because apparently you haven't been learning from previous problems. this is the 3rd or 4th problem with the same type of conditions...

OpenStudy (anonymous):

it doesnt help me...and i work them out and get totally different anwser then from what you get.

OpenStudy (anonymous):

just compare examples in the book that give you point slope form, and look at how the graph is constructed. 2 points make a line right?

OpenStudy (anonymous):

ok, write out what you have... we'll see if they're right...

OpenStudy (anonymous):

x= 2 is all i get...so i guess it would be x=2-y?

OpenStudy (anonymous):

no, x=2 is your answer...

OpenStudy (anonymous):

well then idk what im doing wrong

OpenStudy (anonymous):

whenever you have undefined slope, that is a vertical line. there is no y in your equation.

OpenStudy (anonymous):

so then is just 2

OpenStudy (anonymous):

think about it, there is no way you can get slope from a vertical line. RISE over RUN? theres no run for a line that just goes straight up, in fact it is 0 RUN. what happens when you divide by 0? you cant

OpenStudy (anonymous):

you need an equation... x=2 is the equation of the vertical line.

OpenStudy (anonymous):

@babacusazn , good point...:)

OpenStudy (anonymous):

so is that it then ? is that the answer? yea you have a o. thats just a o....so

OpenStudy (anonymous):

5x=2? is that the answer?

OpenStudy (anonymous):

draw the point (2,5) the location of that point is at position 2 away from the origin on the x axis and 5 up from the y axis. the only thing that could pass through that point with the aforementioned undefined slope is a vertical line. where does that vertical line pass through?

OpenStudy (anonymous):

that doesnt help me....is the answer 5x=2? because that all im getting i dont understand what you are saying...nor do i understand rise over run....

OpenStudy (anonymous):



OpenStudy (anonymous):

is that where the point is located?

OpenStudy (anonymous):